Suppose the production function of a firm is given as X=0.5L ½ K½ price of labor and capital are given as $5 and $10 respectively, and the firm has cos out lay of$ 600



I. The profit maximizing level of L & K to employ.



II. The MRTSL,K at optimum

1
Expert's answer
2022-03-13T19:00:32-0400

production function, X= "0.5l^{0.5}2k^{0.5}"

MPL="\frac{dx}{dl}" ="0.25l^{0.5}2k^{0.5}"

MPK="\frac{dx}{dk}" =".5l^{0.5}k^{0.5}"

MRTS=MPL/MPK=w/r

w=5

r=10

MRTS = ( "0.25l^{0.5}2k^{0.5}" / ("0.5l^{0.5}k^{0.5}" )

"\frac{k}{l}" = "\frac{1}{2}"

K=2L

Calculating the combination of labor that maximizes the firm's output

C = wL + rK, and K=2L

600=5L + "10\times2l"

600=25L

L=24

K=2L = 48

Combination of labor and capital that maximizes the firm's output is;

w,r)=(24,48)

Maximum output, X= "0.5l^{0.5}2k^{0.5}"

X= "0.5(24)^{0.5}2(48)^{0.5}"

=33.96

Maximum output= 33.96